screen is separate from double slit source at 2.5 m the distance between the two slits is 0.05mm the second order bright fringe (m=5)is 9.5cm from the center line .(a) determine the wavelength of the light.(b) calculate the distance between adjaseet bright fringe




1
Expert's answer
2021-04-22T09:30:43-0400


Given

Screen is separated from double slit

D = 2.5m


Distance between two slits ( d) is 0.05 mm


Second order bight fringe is 9.5 cm from center line



Usimh

"\boxed{ y_2= {2\lambda D\over d}}"


"\implies"


"\boxed{ \lambda= {y_2 d\over 2D}}"


"\boxed{ \lambda= {95\times 5 \times 10^{-3}\over 5}}"


"\implies"

(a)

"\boxed { \lambda = 950 nm }"



Distance between adjacent bright fringe

(b)


"\boxed{\Delta y= {\lambda D\over d} = 47.5\times 10^{-3}mm}"